y=2x-4 Plot the y-intercept and find other points using the slope. Connect the points.

Answer by midwood_trail(310)

y = 2x-4 Plot the y-intercept and find other points using the slope. Connect the points.
Select different values for x and simplify.
Let x = -1, 0, and 1 and simplify. You can pick any number for x and simplify but I always choose simple numbers.
Let x = -1
y = 2(-1) - 4
y = -2 - 4
y = -6
Our first point is (-1, -6).
===========================
Let x = 0
y = 2(0) - 4
y = 0 - 4
y = -4
Our second point is (0, -4).
============================
Let x = 1
y = 2(1) - 4
y = 2 - 4
y = -2
Our third point is (1, -2).
Now, get some graph paper and plot the 3 points I just found.
After plotting each point, connect the points with a straight line.
Did you follow?